OpenStudy (anonymous):

The vertex of the function’s graph is given. Find c. y = 2x2 - 6x + c; 3/2, 0.5) a. 5 c. -3 b. -2 d. 2

OpenStudy (anonymous):

x = 3/2 and y = 1/2 so plug them into your equation and solve for c.

OpenStudy (anonymous):

that is what I need help with.

OpenStudy (anonymous):

\(\frac{1}{2}=2\left(\frac{3}{2}\right)^2-6\left(\frac{3}{2}\right)+c\) so what part do you need help with?